OpenStudy (anonymous):

Find the coordinates of the center and the length of the radius of the following circle. (x + 2) 2 + y 2 = 10

OpenStudy (anonymous):

comparing with standard form \[(x-a)^2+(y-b)^2=r^2\] centre is (-2,0) radius=\[\sqrt{10}=3.16\]

OpenStudy (anonymous):

thank you! This is difficult for me.

OpenStudy (anonymous):

write your equation as \[(x-(-2))+(y-0)^2=(\sqrt{10})^2\] now comparing with standard form (-2,0) radius=sqrt(10)=3.16 i hope it is clear